Circular No. 6443

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA SAGITTARII 1996 S. Nakano, Sumoto, Japan, reports the discovery of an apparent nova

by Yukio Sakurai (Mito, Ibaragi-ken) on a Fuji G400 film taken on July 11.590 UT using a 300-mm f/2.8 camera lens; the star was present at mag 10 on both July 11.6 and 12.6. T. Kojima (YGCO Chiyoda Observatory) measured the nova’s position and magnitude on a CCD image taken with a 0.25-m f/6.0 reflector on July 12.673: α = 18^{h}23^{m}42^{s}.49, δ = −18^{o}07′15′′.0 (equinox 2000.0), R = 10.1.

R. M. Wagner and P. Martini, Ohio State University; S. G. Starrfield, Arizona State University; and G. D. Schmidt, University of Arizona, obtained a spectrum (range 320–670 nm, resolution 0.35 nm) of the nova at the Multiple Mirror Telescope on Aug. 4.221 UT, showing emission lines of the Balmer series of hydrogen, Fe II, and Ca II, as well as others, all exhibiting P-Cyg profiles. While the spectrum has not yet been reduced for velocities and other numbers, it appears typical of a relatively slow nova in its early decline phase. The MMT encoders yielded the following position end figures for the nova (cf. Kojima’s position, above): 42^{s}.17, 12′′.6.

A. Gilmore and P. Kilmartin, Mount John University Observatory, report the following photometry of the nova obtained on July 13.498: V = 10.59, U −B = +0.20, B−V = +1.10, V −R = +0.80, V −I = +1.55. They remark that the U count was low but fairly consistent, and the colors look more like those of a reddened nova than a normal star. For comparison, a star ∼ 1′ north of the nova had V = 10.63, U − B = +0.62, B − V = +1.02, V − R = +0.58, V − I = +1.14.

COMET C/1996 J1 (EVANS-DRINKWATER) Ephemeris from orbital elements on MPC 27428:

1996 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
July 26 11^{h}41.^{m}27 −18^{o}53.^{′}6 2.861 2.546 62.^{o}0 20.^{o}6 12.8
Aug. 5 11 55.56 −19 16.8 2.857 2.440 56.0 20.2 12.7
15 12 11.54 −19 51.6 2.843 2.333 50.5 19.6 12.4
25 12 29.23 −20 36.6 2.821 2.228 45.2 18.8 12.2
Sept. 4 12 48.71 −21 30.0 2.788 2.123 40.4 17.9 12.0
14 13 10.10 −22 29.3 2.748 2.020 35.8 16.9 11.7
24 13 33.53 −23 31.9 2.700 1.918 31.5 15.9 11.5
1996 August 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green